OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form?

OpenStudy (anonymous):

x + y = 6 / x – y = 6 / x – y = –6 / x + y = –6

OpenStudy (anonymous):

x – y = 6